OpenStudy (anonymous):

The \(\color{green}{Average}\) of Five numbers is \(\color{blue}{n}\).. If we include two more numbers, then the average becomes: a) Increase by 2. b) Decrease by 2. c) Increase by 1. d) Decrease by 1.

hartnn (hartnn):

are they 5 consecutive numbers?

OpenStudy (anonymous):

Oh sorry, yes they are consecutive numbers..

OpenStudy (anonymous):

(╯° Д° )╯︵sǝʎǝuıɹǝʇɐʍ

OpenStudy (anonymous):

(c)

hartnn (hartnn):

right,so sum =5n say x+x+1+x+2+x+3+x+4=5n---->5x+10=5n or n=x+2 now x+x+1+x+2+x+3+x+4+x+5+x+6=7x+21=7(n-2)+21 average =n-2+3=n+1, so yes it increases by 1

OpenStudy (anonymous):

Yes mukushla you are right..

OpenStudy (anonymous):

Well Done @hartnn

hartnn (hartnn):

thanks :)

OpenStudy (anonymous):

do u like the new pic..my profile guys?

OpenStudy (anonymous):

It is very nice.. I wish I could give medal to your profile pic too.. ha ha ha..

hartnn (hartnn):

are they underwater? seems so romantic.

OpenStudy (anonymous):

this is from an animation that i love it sooooooo much

hartnn (hartnn):

whats the name of that anime?

OpenStudy (anonymous):

which anime?

OpenStudy (anonymous):

Spirited Away

OpenStudy (anonymous):

thats my number 1 anime ever

OpenStudy (anonymous):

@nphuongsun93 how did you write that ??

hartnn (hartnn):

is that a 2001 movie?spirited away?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

(╯° Д° )╯︵┻━┻ copy paste off the internet ǝuoʎɹǝʌǝ ɐןɥsnʞnɯ sǝʎǝuıɹǝʇɐʍ http://www.fliptext.org/

hartnn (hartnn):

36unsbuonɥdu

OpenStudy (anonymous):

ʎɐʍɐ pǝʇıɹıds

OpenStudy (anonymous):

thats nice @nphuongsun93

OpenStudy (anonymous):

(╯° Д° )╯︵ uuʇɹɐɥ waterineyes ノ(º_ºノ)

hartnn (hartnn):

@nphuongsun93 if your name did not contain p,then even after flipping text,all letters make sense....

hartnn (hartnn):

sorry h

OpenStudy (anonymous):

ಠ_ಠ (・ω・) not "3" :P

OpenStudy (anonymous):

well imma download spirited away l0l

hartnn (hartnn):

i already downloaded the 3.14 Gb print of spirited away,and already started watching....

OpenStudy (anonymous):

u kidding

OpenStudy (anonymous):

lol

hartnn (hartnn):

not kidding at all!!

OpenStudy (anonymous):

man what is ur internet speed there it must be 3-4 Mb/sec or something

hartnn (hartnn):

i downloaded it from LAN here,with slowest speed of 11Mbps....

OpenStudy (anonymous):

o.O o.O O.o o.O o.O o,O

OpenStudy (anonymous):

=.= 11mbps.. downloading from torrent 150kb ps ._.

hartnn (hartnn):

and your pic is is from the scene after 1:50 hrs from beginning

OpenStudy (anonymous):

700mb

OpenStudy (experimentx):

had watched that couple of months ago ... it was good

OpenStudy (anonymous):

man what is ur number one anim?

OpenStudy (anonymous):

all of u?

OpenStudy (anonymous):

code geass

OpenStudy (experimentx):

that was nice too ..

OpenStudy (anonymous):

man thats 8.9 in imdb but not in the top 250...how its possible?

OpenStudy (experimentx):

can you show the list?

OpenStudy (anonymous):

http://www.imdb.com/chart/top

OpenStudy (experimentx):

44th

hartnn (hartnn):

i have seen an anime TV series,Death note,it was awesome....

OpenStudy (experimentx):

Death note is intriguging

OpenStudy (experimentx):

from movie type genre ... my favouraite http://en.wikipedia.org/wiki/Grave_of_the_Fireflies

OpenStudy (experimentx):

last anime i watched was Cowboy Bebop ... and it was very nice too

OpenStudy (anonymous):

Ha ha ha ha... And My Internet speed is far better than you all.. Can anyone guess ??

OpenStudy (experimentx):

4 mbps?

OpenStudy (anonymous):

I can give you one more chance as you are my best buddy too like @mukushla 4 Mbps is not better than @hartnn 's 11 Mbps..

OpenStudy (experimentx):

if I copy files from C: to D: the speed is 24 mbps

hartnn (hartnn):

i said slowest is 11mbps,highest was 70 mbps!! do u have better than 70Mbps????

OpenStudy (anonymous):

@mukushla I think I can understand you and your internet speed.. ha ha ha ha..

OpenStudy (experimentx):

you guys have crazy speed

OpenStudy (anonymous):

Wanna know my internet speed ?? It is: 30 Kbps.. And now I am serious..

OpenStudy (anonymous):

hahahaha....... @waterineyes man im gonna blow here.hahhhaaaaaaaaahaaaaaa

OpenStudy (anonymous):

wait is this internet speed or wud ´Д` my internet's min = 30kb ps, max = 300kb ps barely enough to surf internet ;-;

OpenStudy (anonymous):

thats exactly what i have here

OpenStudy (anonymous):

ha ha ha ha ha ha ha ha ha.. @mukushla That is why I said I understand you.. Kazem..

OpenStudy (experimentx):

i don't have good speed here either ... 8 yrs back i remember surfing on 4 kbps

OpenStudy (anonymous):

8 years is very long time @experimentX

OpenStudy (experimentx):

probably 10 yrs back ... in 2002

OpenStudy (anonymous):

Okay guys, now it is time to go to college and attend the boring lectures.. I have to go now.. Miss you guys.. Bye...

OpenStudy (experimentx):

have fun ...

hartnn (hartnn):

that was my LAN speed,the net speed i around 1Mbps,and with IDM it once reached 5Mbps!

OpenStudy (anonymous):

Sure but I have fun on OS more instead of college..

OpenStudy (anonymous):

@hartnn thats superinternet or something u have man

OpenStudy (experimentx):

yeah ... i agree that. since i skipped all my classes and here I'm now .... 6 days before exam :((

OpenStudy (anonymous):

@hartnn is not living on Earth he or she is living in Heaven..

OpenStudy (experimentx):

that's pretty crazy speed ... min: 23 kbps max 150 kbps ... barely able to stream youtube.

OpenStudy (anonymous):

I can't even open a separate Tab while I am using OS.. That is my condition on this speed..

OpenStudy (experimentx):

probably this works with comet + ajax ... shouldn't take much bandwidth.

hartnn (hartnn):

we have own internet network of our college with our own infrastructure, such net speeds are accessible to every1 staying here in hostels, which helps them to stream seamlessly any video lectures from the entire net......instead we stream live matches :P

OpenStudy (anonymous):

Ok Finally I leave.. Bye..

OpenStudy (experimentx):

you should change your browser ... and tweak up your pc

OpenStudy (experimentx):

never seen a girl who likes DBZ

OpenStudy (anonymous):

who?

hartnn (hartnn):

girls don't like action much

OpenStudy (experimentx):

anyone interested in PDE

OpenStudy (experimentx):

yeah ... kinda ... had a friend who was crazy DB fan.

hartnn (hartnn):

whats PDE, a form of Differential Equation?

OpenStudy (experimentx):

partial DE